A researcher dissolves 2.45g MgCl2 of 250g of water. Find the boiling point elevation elevation for this solution assuming MgCl2 dissociates completely and Kb = 0.512° C/m for H2O show all work
T = Kbm
1
Expert's answer
2016-02-16T08:57:02-0500
Molar mass of PCl is 66.4268 g/mol
53,9 g/ 66.4268 g/mol = 0,81 mol
A mole is a measurement that represents 6.02 * 10^23 particles
0,81 mol * 6.02 * 10^23 particles = 4,87 * 10^23 particles
In 53.9 grams of PCl is 4,87 x 10^23 particles